1. Use suitable identims to find the following products:
(i) (% + 4)(:c + 10) (ii) (:c + 8)(:c - 10)

(iii) (S:c + 4)(3x -•SJ
(v) (3 - 2:c)(8 + 2%).

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
(i) {:c + 4)(.:c + 10) = x2 + (4 + 10>.t + (4 x 10) [Using identity: (:r + a)(:r + b) = x2 + (a + b)x + ab] = x2 + 14% + 40. 37 (ii (% + 8)(% - 10) = x2 + (8 - O)x + 8(- 10) (Using identity: (% + a)(x + b) = x2 + (a + b')x + ab] = x2 - 2x - 80. (iii (3% +4)(& -5) = (3%>2 + (4 - 5X&> + 4 x < 5) [Using identity: (.x + a)(.x + b) = x2 + (a + b")% + ab] = 9x2 - 3% - 20. (iv) (y2 +i)(y2 -i) = < >2 - (iJ (Using identity:(.x + a)(x - a) = x2 - a2J = y - 4 • (v) (3 - 2x)(3 + 2x) = (3)2 - (2x)2 [Using identity: (x + a)(.r - a) = x2 - a2J = 9 - 4x2.